OpenStudy (anonymous):

Use the substitution method to solve the linear system. 3x=9 -2x+y=-8 and(different problem) x-2y=-13 y=-2x-6

OpenStudy (anonymous):

y=-2 and x=3 ^ for the first one

OpenStudy (anonymous):

x+4x+12=-13 5x=-25 x=-5 y=-2(-5)-6 y=10-6 y=4 x=-5 and y=4 for the second one :)